solve by substitution method
6x+7y=14
x= 25-4y

Answer by mananth(12270)

6 x + 7 y = 14 .............1
4 y = 25 -1 x ---------2
/ 4 =
y = 6.25 -0.25 x
Substitute y in (1)
6 x + 7 ( 6.25 -0.25 x ) = 14
6 x + 43.75 + -1.75 x = 14
6 x -1.75 x = 14 -43.75
4.25 x = -29.75
/ 4.25
x= -7
Plug the value of x in (1)
6 x + 7 y = 14
6 * -7 + 7 y = 14
-42 + 7 y = 14
7 y = 56
/ 7
y= 8
m.ananth@hotmail.ca